OpenStudy (anonymous):

Solve. 9x - 5 is less than or equal to 7x - 11

OpenStudy (anonymous):

\[9x - 5 \le 7x - 11\]

OpenStudy (anonymous):

i got \[x \le -3\], is that right?

OpenStudy (jhannybean):

\[9x-7x \le -11 +5\]\[2x\le -6\]\[x\le ~?\]

OpenStudy (dinamix):

yup its @highschoolstuent55236 u are good ;D

OpenStudy (anonymous):

okay thanks